5 years ago Mary was 5 times as old as her Son. 5 years hence her age will be 8 less than three times the corresponding age of her Son. Find the age of Mary and her Son
Answer by mananth(16949)

5 years ago Mary was 5 times as old as her Son. 5 years hence her age will be 8 less than three times the corresponding age of her Son. Find the age of Mary and her Son
Let son's age be x now
let Mary'e age be y now
5 years ago
y-5=5(x-5)
y-5= 5x-25
y-5x=-20
.
5 years later
son's age will be x+5
Mary's age will be y+8
y+8 = 3(x+8)-8
y+8= 3x+24 -8
y-3x=8
y-5x-y+3x= -20-8
-2x=-28
x=14 son's age now
plg the value of x in y-3x=8
y-42=8
y= 50 mother's age now